# Scalable Oversight for Superhuman AI via Recursive Self-Critiquing

### v3

@@ Appendix B Prompts for AI-AI Experiments
We adopt the following prompt template in Figure 4, 5, 6, 7 to conduct response generation and multi-stage critiques. Additionally, our smaller SFT models, particularly those with 0.5B parameters and limited capabilities, occasionally fail to follow instructions properly. To address this issue, we incorporate hints in the output section to enhance the model’s instruction adherence and chain-of-thought analysis process. We set the sampling temperature to 1.0 and top-p to 1.0.
@@
Appendix C Supplemental AI Recursive Self-critiquing Experiment
In this seciton, we investigate the effectiveness of the recursive self-critiquing across differernt LLMs.
C.1 Setup
We utilize multiple reasoning, knowledge, and alignment-related datasets, including MATH, GPQA, TruthfulQA, MMLU-Pro, and BoolQ. Below is a detailed introduction to each task category.
  • MATH[Hendrycks et al., 2021] is a mathematical problem-solving dataset consisting of 12,500 challenging competition-level math problems, designed to assess machine learning models’ mathematical reasoning abilities. Each problem is accompanied by a fully worked-out step-by-step solution, enabling models to learn how to generate answer derivations and explanations.
  • GPQA[Rein et al., 2023] is a highly challenging multiple-choice question dataset consisting of 448 questions crafted by domain experts in biology, physics, and chemistry. The dataset is designed to assess the reasoning capabilities of both human experts and state-of-the-art AI models on complex scientific topics. To ensure its difficulty and quality, questions were validated by experts with PhD-level knowledge, achieving an accuracy of only 65% (or 74% after correcting clear retrospective mistakes). In contrast, highly skilled non-expert validators, even with unrestricted web access for over 30 minutes per question, achieved only 34% accuracy.
  • TruthfulQA[Lin et al., 2022] evaluates the truthfulness of language models in answering questions, comprising 817 questions across 38 categories, including health, law, finance, and politics. The questions were carefully designed to reflect common human misconceptions or false beliefs, making them particularly challenging. To perform well, models must avoid generating false answers learned from imitating human-written text, which often contains misinformation.
  • BoolQ[Clark et al., 2019] is a reading comprehension dataset designed to study naturally occurring yes/no questions, meaning questions that arise spontaneously in unprompted and unconstrained settings. The dataset presents unexpected challenges, as its questions often involve complex, non-factoid information and require entailment-like inference rather than simple fact retrieval.
  • MMLU-Pro[Wang et al., 2024b] is an enhanced version of MMLU designed to go beyond MMLU’s primarily knowledge-driven evaluation. MMLU-Pro incorporates more challenging reasoning-focused questions, expands the answer choice set from 4 to 10 options, and removes trivial and noisy questions from MMLU. Experimental results show that MMLU-Pro significantly increases difficulty, leading to an accuracy drop of 16% to 33% compared to MMLU.
Figure 8: The Sampling Strategy of AI Self Recursive Critiquing.
We employ the structured prompt illustrated in Figure 10, 11, 12, 13, 14, 15 to obtain consistent forms of response, critique, and higher-order critique across different models and datasets. Given the variations in how different models adhere to and comprehend instructions, the prompt structure is slightly adjusted for each model. These adjustments primarily focus on constraints related to output length and the format of the decision-making answers.
We adopt consistent metrics and baselines as in human experiments. Each score in the experiments is averaged over 10 different runs. To ensure fairness across different stages of effort, we followed the sampling strategy illustrated in Figure 8 to sample the responses of the model to a question and critics of various orders. Each sampling begins by obtaining 7 responses to the same question. From the first two responses, we further derive 5 critics. Similarly, we generate 3 critics of critics and 1 critics of critics of critics. To ensure the reliability of the results, we repeat the entire process 10 times for the same question and report the average outcomes of these ten iterations. To enhance the diversity of the sampling process, we set the sampling temperature to 1.0 and top-p to 0.95.
C.2 Experimental Results
Figure 9: The relative accuracy improvement of critique and recursive critique stages compared to the response stage. Scores are averaged across all datasets. The improvement is calculated as exp(Acc_stage - Acc_response), where samples are grouped according to their response accuracy levels.
Current AI Models struggle to perform recursive critique tasks.
We plot the average improvement of critique and recursive cirtics over response stage in Figure 9, where we find that the performance of critiques and higher-order critiques remains constrained in most scenarios. In detail, critique, critique of critique ($C^{2}$) and third-order critiques ($C^{3}$) struggle to surpass the accuracy levels achieved at the Response stages. This phenomenon suggests that current models may have limited capacity for self-critiquing, which is also partially validated in the work of Huang et al. [2023],  Tang et al. [2025] and summarized by Kamoi et al. [2024]. This finding further highlight the importance of post-training models to perform better critique tasks as in McAleese et al. [2024]. Tables 13 and 13 include the results from Gemma2 9B Instruct and Qwen-2.5 14B Instruct on MMLU-Pro and BoolQ.
Table 9: Performance comparison of AI self recursive critique using Gemma2 9B Instruct. We select the question set that $Q^{\prime}=\{q\mid 0<\text{Acc}(q)<0.7,q\in Q\}$.
Dataset Stage Accuracy Majority Voting Naive Voting
MATH Response 22.82 19.90 22.53
Critic 26.14 25.23 23.30
$C^{2}$ 26.90 26.60 25.00
$C^{3}$ 27.32 27.32 25.69
GPQA Response 19.68 16.24 19.76
Critic 24.43 23.92 19.57
$C^{2}$ 22.60 22.31 20.39
$C^{3}$ 22.63 22.63 20.75
TruthfulQA Response 24.74 22.63 23.16
Critic 39.98 39.37 29.68
$C^{2}$ 34.67 35.68 30.74
$C^{3}$ 37.26 37.26 32.11
Table 10: Performance comparison of AI self recursive critiquing using Qwen2.5 14B Instruct. We select the question set that $Q^{\prime}=\{q\mid 0<\text{Acc}(q)<0.7,q\in Q\}$.
Dataset Stage Accuracy Majority Voting Naive Voting
MATH Response 31.69 31.19 30.76
Critic 34.56 34.81 34.27
$C^{2}$ 35.19 34.92 34.86
$C^{3}$ 35.89 35.89 35.41
GPQA Response 22.09 19.56 21.69
Critic 23.84 23.46 23.05
$C^{2}$ 23.30 23.24 22.50
$C^{3}$ 24.26 24.26 23.35
TruthfulQA Response 25.73 22.37 27.32
Critic 39.57 38.45 34.12
$C^{2}$ 37.87 37.84 34.54
$C^{3}$ 38.66 38.66 36.49
Table 11: Performance comparison between Gemma2 9B and Qwen2.5 14B models, where C denotes correct and W denotes wrong. For example, 1C 1W means one correct response and one wrong response were input to the critic stage.
Input Type Gemma2 9B Qwen2.5 14B
Critic Accuracy 1C 1W 42.3% 55.5%
2C 64.3% 98.4%
2W 13.6% 1.1%
$C^{2}$ Accuracy 1C 1W 46.5% 55.7%
2C 89.8% 97.1%
2W 4.8% 1.6%
$C^{3}$ Accuracy 1C 1W 51.1% 52.3%
2C 92.8% 98.9%
2W 2.7% 1.3%
Potential effectiveness in specific models.
We further compare the performance of Qwen and Gemma model across different datasets, following the setting in Section 4.2. The results are shown in Table 10 and Table 10. From these results, we can observe that the disparity of higher-order critiquing ability exists in specific models. Although performance is limited, Qwen-2.5-14B-Instruct exhibits more effective in high-order critiquing, as we can observe progressive improvements between response and recursive critiques across stages. Further performance gap analysis in Table 11 reveals that the main discrepancies may arise from those gap in distinguishing the true statements from mixed true and false inputs.
Table 12: Performance comparison of AI self recursive critiquing using Gemma2 9B Instruct. We select the question set that $Q^{\prime}=\{q\mid 0<\text{Acc}(q)<0.7,q\in Q\}$.
Dataset Stage Accuracy Majority Voting Naive Voting
MMLU Pro Response 22.31 25.43 25.84
Critic 32.95 32.81 28.90
$C^{2}$ 32.25 32.24 30.35
$C^{3}$ 31.79 31.79 31.04
BoolQ Response 31.36 28.78 33.66
Critic 32.59 30.24 31.22
$C^{2}$ 29.67 28.05 27.80
$C^{3}$ 32.44 32.44 27.07
Table 13: Performance comparison of AI self recursive critiquing using Qwen2.5 14B Instruct. We select the question set that $Q^{\prime}=\{q\mid 0<\text{Acc}(q)<0.7,q\in Q\}$.
Dataset Stage Accuracy Majority Voting Naive Voting
MMLU Pro Response 34.71 35.58 33.75
Critic 35.50 35.58 35.17
$C^{2}$ 35.78 35.67 35.42
$C^{3}$ 36.83 36.83 35.25
BoolQ Response 25.98 20.41 27.35
Critic 27.14 24.49 27.35
$C^{2}$ 26.53 26.12 25.92
$C^{3}$ 28.16 28.16 25.51
@@ C.2 Experimental Results
Appendix D Broader Impacts
Our recursive self-critiquing framework offers significant potential for maintaining effective AI oversight as capabilities surpass human abilities, potentially democratizing participation in AI governance through more tractable evaluation processes. However, this approach also carries risks, including the possibility of creating false confidence in oversight effectiveness given current limitations in AI critique capabilities, vulnerability to adversarial examples designed to bypass critique chains, and concerns about computational resource requirements potentially limiting access to effective oversight tools among less-resourced stakeholders. We acknowledge these potential impacts and encourage continued research to strengthen the reliability of scalable oversight methods.

### v4

@@ Appendix B Prompts for AI-AI Experiments
We adopt the following prompt template in Figure 5, 6, 7, 8 to conduct response generation and multi-stage critiques. Additionally, our smaller SFT models, particularly those with 0.5B parameters and limited capabilities, occasionally fail to follow instructions properly. To address this issue, we incorporate hints in the output section to enhance the model’s instruction adherence and chain-of-thought analysis process. We set the sampling temperature to 1.0 and top_p to 1.0.
Appendix C Recursive Self-critiquing on Differernt LLMs
In this section, we further investigate the effectiveness of recursive self-critiquing across different LLMs on various tasks.
Table 4: Performance comparison of AI self recursive critiquing. We select the question set that $Q^{\prime}=\{q\mid 0<\text{Acc}(q)<0.7,q\in Q\}$ to focus on questions where initial accuracy is moderate, as questions with very high initial accuracy leave limited room for meaningful improvement through recursive self-critiquing.
Dataset Stage Qwen2.5-14B-Instruct Gemma2-9B-Instruct Qwen3-30B-A3B-Thinking-2507
Accuracy Majority Naive Accuracy Majority Naive Accuracy Majority Naive
MMLU Pro Response 34.71 35.58 22.31 25.43 37.21 36.80
Critic 35.50 35.58 35.17 32.95 32.81 28.90 41.62 41.98 40.00
$C^{2}$ 35.78 35.67 35.42 32.25 32.24 30.35 40.81 41.09 40.91
$C^{3}$ 36.83 35.25 31.79 31.04 37.15 37.56
BoolQ Response 25.98 20.41 31.36 28.78 33.40 28.09
Critic 27.14 24.49 27.35 32.59 30.24 31.22 33.70 32.68 32.15
$C^{2}$ 26.53 26.12 25.92 29.67 28.05 27.80 32.20 32.06 32.15
$C^{3}$ 28.16 25.51 32.44 27.07 34.03 32.87
MATH Response 31.69 31.19 22.82 19.90 52.56 61.86
Critic 34.56 34.81 34.27 26.14 25.23 23.30 50.85 52.27 52.65
$C^{2}$ 35.19 34.92 34.86 26.90 26.60 25.00 41.76 42.61 43.18
$C^{3}$ 35.89 35.41 27.32 25.69 36.08 36.36
GPQA Response 22.09 19.56 19.68 16.24 38.14 40.89
Critic 23.84 23.46 23.05 24.43 23.92 19.57 38.82 39.88 40.44
$C^{2}$ 23.30 23.24 22.50 22.60 22.31 20.39 37.74 37.55 37.34
$C^{3}$ 24.26 23.35 22.63 20.75 35.70 35.68
TruthfulQA Response 25.73 22.37 24.74 22.63 36.21 36.26
Critic 39.57 38.45 34.12 39.98 39.37 29.68 43.97 44.06 43.30
$C^{2}$ 37.87 37.84 34.54 34.67 35.68 30.74 44.18 44.25 43.97
$C^{3}$ 38.66 36.49 37.26 32.11 44.47 44.25
C.1 Datasets
We utilize reasoning, knowledge, and alignment-related datasets, including the following:
  • MATH(Hendrycks et al., 2021) is a mathematical problem-solving dataset consisting of 12,500 challenging competition-level math problems, designed to assess machine learning models’ mathematical reasoning abilities. Each problem is accompanied by a fully worked-out step-by-step solution, enabling models to learn how to generate answer derivations and explanations.
  • GPQA(Rein et al., 2023) is a highly challenging multiple-choice question dataset consisting of 448 questions crafted by domain experts in biology, physics, and chemistry. The dataset is designed to assess the reasoning capabilities of both human experts and state-of-the-art AI models on complex scientific topics. To ensure its difficulty and quality, questions were validated by experts with PhD-level knowledge, achieving an accuracy of only 65% (or 74% after correcting clear retrospective mistakes). In contrast, highly skilled non-expert validators, even with unrestricted web access for over 30 minutes per question, achieved only 34% accuracy.
  • TruthfulQA(Lin et al., 2022) evaluates the truthfulness of language models in answering questions, comprising 817 questions across 38 categories, including health, law, finance, and politics. The questions were carefully designed to reflect common human misconceptions or false beliefs, making them particularly challenging. To perform well, models must avoid generating false answers learned from imitating human-written text, which often contains misinformation.
  • BoolQ(Clark et al., 2019) is a reading comprehension dataset designed to study naturally occurring yes/no questions, meaning questions that arise spontaneously in unprompted and unconstrained settings. The dataset presents unexpected challenges, as its questions often involve complex, non-factoid information and require entailment-like inference rather than simple fact retrieval.
  • MMLU-Pro(Wang et al., 2024b) is an enhanced version of MMLU designed to go beyond MMLU’s primarily knowledge-driven evaluation. MMLU-Pro incorporates more challenging reasoning-focused questions, expands the answer choice set from 4 to 10 options, and removes trivial and noisy questions from MMLU. Experimental results show that MMLU-Pro significantly increases difficulty, leading to an accuracy drop of 16% to 33% compared to MMLU.
C.2 Setup
Models and Configuration
We conduct experiments with three state-of-the-art language models: Qwen2.5-14B-Instruct (Qwen et al., 2025), Gemma2-9B-Instruct (Team et al., ), and Qwen3-30B-A3B-Thinking-2507 (Team, 2025). For Qwen2.5-14B-Instruct and Gemma2-9B-Instruct, we set temperature=1.0 and top-p=0.95 to enhance sampling diversity. For Qwen3-30B-A3B-Thinking-2507, we follow official recommended settings with temperature=0.6, top-p=0.95, top-k=20, and min-p=0, extracting only the response content while excluding the thinking traces. We employ structured prompts to ensure consistent response formats across models and datasets (Figures 9, 10, 11, 12, 13, and 14), with minor model-specific adjustments for output length constraints and answer formatting.
Sampling Strategy and Evaluation
To ensure fair comparison across recursive critique stages, we follow a controlled hierarchical sampling strategy. For Qwen2.5-14B-Instruct and Gemma2-9B-Instruct, we generate 7 initial responses per question, derive 5 critics from the first two responses, then produce 3 critics of critics ($C^{2}$) and 1 critic of critics of critics ($C^{3}$). This process repeats 10 times per question with averaged results. For Qwen3-30B-A3B-Thinking-2507, due to longer output length and resource constraints, we use 16 initial responses with 4 critic pairs each producing 4 critiques at each subsequent level ($C^{2}$ and $C^{3}$). This pyramidal structure balances computational efficiency with evaluation diversity. We adopt identical evaluation metrics and baselines as used in the human experiments.
Table 5: Performance comparison by initial response correctness. C denotes correct, W denotes wrong responses. 1C1W/2C/2W indicate different combinations of initial response correctness. AI results from MATH dataset, human results from multiple datasets.
(a) AI performance
Stage Type Gemma2 Qwen2.5
Critic 1C1W 42.3 55.5
2C 64.3 98.4
2W 13.6 1.1
$C^{2}$ 1C1W 46.5 55.7
2C 89.8 97.1
2W 4.8 1.6
$C^{3}$ 1C1W 51.1 52.3
2C 92.8 98.9
2W 2.7 1.3
(b) Human performance
Stage Type Accuracy
Critic 1C1W 56.6
2C 86.8
2W 21.0
$C^{2}$ 1C1W 73.7
2C 93.8
2W 29.0
$C^{3}$ 1C1W 75.9
2C 93.5
2W 31.2
C.3 Experimental Results
Potential effectiveness in specific models.
The results in Table 4 compare the performance of Qwen and Gemma models across different datasets. From these results, we observe clear disparities in higher-order critiquing abilities across both models and benchmarks. Qwen2.5-14B-Instruct demonstrates stronger recursive-critiquing effectiveness, exhibiting consistent improvements from the initial response to deeper critique stages. However, other models, such as Gemma2-9B-Instruct or Qwen3-30B-A3B-Thinking-2507, show gains only on a limited subset of datasets.
These performance gaps likely stem from difficulties in distinguishing true statements from inputs containing mixed true–false information, as illustrated in Table 5. For comparison, Table 5 reports the accuracy of our human annotators across different input conditions. Human performance not only far exceeds that of AI models in both the 1C1W and 2W scenarios, but also shows substantial improvement as the length of the recursive-critique chain increases.
Current AI models show limited capabiliy in self-recursive critique.
We further investigate recursive self-critique performance across different large models and accuracy intervals. Testing models ranging from Qwen2.5-7B to 72B, we find that models typically demonstrate self-critique effectiveness in intervals where response accuracy is relatively moderate. However, overall we observe that models’ self-critique capabilities are limited, with typically modest improvement margins. These results are also partially validated in prior work (Huang et al., 2023; Tang et al., 2025) and summarized by Kamoi et al. (2024). This finding further highlights the importance of investigating approaches to improve models’ critique performance (McAleese et al., 2024).
Figure 4: The relative accuracy improvement of critique and recursive critique stages compared to the response stage. Scores are averaged across all datasets. The improvement is calculated as $\exp(\text{Acc}_{\text{stage}}-\text{Acc}_{\text{response}})$, where samples are grouped according to their response accuracy levels.
Nevertheless, we note that these limitations do not diminish the potential of recursive self-critiquing as a scalable oversight paradigm. A lthough current models’ self-critique abilities require improvement, recursive self-critiquing can still yield improvements in weak-to-strong settings as demonstrated in Section 5. This aligns with scalable oversight scenarios where AI provides effective supervision signals when superior to humans.
@@ C.3 Experimental Results
Table 6: High quality and low quality response examples.
Quality Definition Type Example and Translation
Math Origin: $f(x)=2\cos^{2}x-\sin^{2}x+2$$\cos 2x=2\cos^{2}x-1$$2\cos^{2}x=\cos 2x+1$$\sin^{2}x+\cos^{2}x=1$$\sin^{2}x=(1-\cos 2x)/2$$f(x)=\frac{3}{2}\cos 2x+\frac{5}{2}$。通过周期计算和最值分析,得到答案B。 Translated: First simplify $f(x)=2\cos^{2}x-\sin^{2}x+2$. Using double angle formula $\cos 2x=2\cos^{2}x-1$, we get $2\cos^{2}x=\cos 2x+1$. Since $\sin^{2}x+\cos^{2}x=1$, we derive $\sin^{2}x=(1-\cos 2x)/2$. Finally $f(x)=\frac{3}{2}\cos 2x+\frac{5}{2}$. Through period calculation and maximum analysis, we arrive at answer B.
[…]
Table 7: High quality and low quality critique annotation examples.
Question Response1 Response2
High Quality Examples
Low Quality Examples
认可response 1的答案,不同意Response2的答案,答案应该是(B)。
Question Response1 Response2
High Quality Examples
Low Quality Examples
Question Response 1 Response 2
Critic 1 Critic 2
High Quality Examples
Low Quality Examples
Critc 1/2的答案是对,应该是(B)。
Question Response 1 Response 2
Critic 1 Critic 2
High Quality Examples
Low Quality Examples
Critics 1/2 are correct, the answer should be (B).